OpenStudy (anonymous):

(x -5)2 + (y + 2)2 = 64 The radius of the circle is ________ units.

OpenStudy (ghazi):

(x-x1)^2+(y-y1)^2=r^2

OpenStudy (ghazi):

now you can calculate the radius

OpenStudy (anonymous):

calculate the radius? ParthKohli could you help?

OpenStudy (ghazi):

just try to reform the given equation as per the general equation (x-5)^2+(y+2)^2=(8)^2....now compare this with general equation

OpenStudy (anonymous):

is the answer 8?

OpenStudy (ghazi):

yes